OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 2x - cos x

OpenStudy (anonymous):

Seems easy! What am I missing?

OpenStudy (campbell_st):

well sin(2x) = 2sin(x)cos(x) substitute this and then perhaps factor the expression

OpenStudy (solomonzelman):

if you need an expression in terms of an angle of a single x, then do what campbell says. use that rule. If you have any "why" questions, or any other question, don't hesitate to ask.

OpenStudy (anonymous):

I just don't understand how to substitute and factor when I'm only given sin2x and -cosx :(

OpenStudy (solomonzelman):

lets do this again. You have: \(\Large\color{black}{ \sin(2x)-\cos(x) }\) and there is a rule, \(\Large\color{blue}{ \sin(\color{red}{a}+\color{magenta}{b})=\sin(\color{red}{a})\cos(b)+\sin(\color{magenta}{b})\cos(\color{red}{a}) }\)

OpenStudy (solomonzelman):

in your situation however, \(\Large\color{black}{ \sin(2x)}\) is same as \(\Large\color{black}{ \sin(x+x)}\) and that means that your a and b are the same, they are both x. this is what you get, \(\Large\color{blue}{ \color{black}{ \sin(2x)}= \sin(\color{red}{x}+\color{magenta}{x})=\sin(\color{red}{x})\cos(x)+\sin(\color{magenta}{x})\cos(\color{red}{x}) }\)

OpenStudy (solomonzelman):

and this means that you have \(\Large\color{blue}{ \sin(\color{red}{x})\cos(x)+\sin(\color{magenta}{x})\cos(\color{red}{x}) }\) which would be \(\Large\color{blue}{ 2\sin(\color{red}{x})\cos(x) }\) just like \(\Large\color{blue}{ C+C=2C }\)

OpenStudy (solomonzelman):

So, so far, do you understand why \(\Large\color{blue}{ \sin(2x)=2\sin(x)\cos(x) }\)

OpenStudy (anonymous):

Yeah, I follow! Thank you for the explanation haha

OpenStudy (solomonzelman):

So, when you have \(\Large\color{blue}{ \sin(2x)-\cos(x) }\) it becomes, \(\Large\color{blue}{2 \sin(x) \cos(x)-\cos(x) }\).